# Unstable entropy for Anosov diffeomorphisms on the $3$-torus

### v1

@@ 2.1.1. Definition
are continuous and satisfy $\lambda_{x}^{s}<1<\lambda_{x}^{c}<\lambda_{x}^{u}$ for any $x\in\mathbb{T}^{3}$. In particular, $f$ is an Anosov map with stable bundle $E_{s}$ and unstable bundle $E_{cu}:=E_{c}\oplus E_{u}$. The bundles $E_{c}$ and $E_{u}$ are respectively called the center and strong unstable bundle of $f$. Finally, $E_{cs}:=E_{c}\oplus E_{s}$ is the center stable bundle of $f$. For a detailed introduction to Anosov diffeomorphisms, we refer to [21].
@@ 2.1.2. Invariant manifolds
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Since $f$ is Anosov, $E_{s}$ and $E_{cu}$ are integrable to $f$-invariant foliations $\mathcal{W}^{s}$ and $\mathcal{W}^{cu}$ called respectively the stable and unstable foliations of $f$, see [21, Theorem 6.2.3]. Moreover, since $E_{u}$ is the strong unstable foliation of a strongly partially hyperbolic diffeomorphism, it integrates to a $f$-invariant foliation $\mathcal{W}^{u}$ called the strong unstable foliation of $f$, see [33, Theorem 4.1]. Actually, by the work of Brin, Burago and Ivanov [6] and Potrie [34], any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is dynamically coherent. This means that the center stable bundle $E_{cs}$ also integrates to a $f$-invariant foliation $\mathcal{W}^{cs}$ called the center stable foliation of $f$. The collection of leaves $\mathcal{W}^{c}(x):=\mathcal{W}^{cs}(x)\cap\mathcal{W}^{cu}(x)$ for $x\in\mathbb{T}^{3}$ defines the center foliation which integrates the center bundle $E_{c}$. For any $\bullet=s,c,u,cs,cu$, any $x\in\mathbb{T}^{3}$ and any $\delta>0$, let $\mathcal{W}^{\bullet}(x,\delta):=\{y\in\mathcal{W}^{\bullet}(x)\mid d^{\bullet}(x,y)<\delta\}$, where $d^{\bullet}$ is the distance on the leaf.
@@ 2.1.3. Local product structure and rectangles
Rectangles exist because of the local product structure, see [12, Eq. $(2.2)$].
@@ 2.1.4. Holonomies
The strong unstable foliation is absolutely continuous, see (5.5) for the exact statement we will need. This fact will be used in the proof of Proposition 5.2 to show the $\mathrm{Hol}^{u}$-invariance of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$. The center stable foliation is not always absolutely continuous and this explains why proving the $\mathrm{Hol}^{cs}$-invariance of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is harder. We refer to [1, §2.2.3] for a more detailed discussion of the lack of regularity of center stable holonomies.
@@ 2.2.1. Conditional measures
We recall some facts about measurable partitions and conditional measures and refer to [22] for more details. Let $(X,\mu)$ be a probability space. For $\xi$ a partition of $X$ into $\mu$-measurable sets, we write $\xi(x)\in\xi$ for the element of the partition containing $x$. Let $\xi_{1},\xi_{2}$ be two partitions. We say that $\xi_{1}$ refines $\xi_{2}$ denoted $\xi_{2}\prec\xi_{1}$ if for $\mu$-a.e. $x\in X$, one has $\xi_{1}(x)\subset\xi_{2}(x)$. The joint of $\xi_{1}$ and $\xi_{2}$ is the partition defined by $\xi_{1}\vee\xi_{2}:=\{\xi_{1}(x)\cap\xi_{2}(x)\mid x\in X\}$. It is the least fine partition that refines both $\xi_{1}$ and $\xi_{2}$. A partition $\xi$ is said to be measurable if there is a sequence of finite partitions $\xi_{n}$ such that $\xi=\vee_{n=0}^{+\infty}\xi_{n}.$
@@
see [12, Equation $(2.6)$]. One defines the conditionals of $\mu$ on the strong unstable manifolds $(\mu^{u}_{x})_{x\in R}$ by replacing $cs$ by $u$ in the previous definition. Although this is not explicit in the notations, the conditionals depend on the rectangle $R$. Note however that the ambiguity only consists in a normalizing constant, see [12, Lemma 2.10].
@@ 2.2.2. Metric unstable entropy
We now recall the definition of the metric unstable entropy. For further details, we refer to [24]. In this subsection, we consider $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and $\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})$ a $f$-invariant probability measure. For a partition $\xi$, define its diameter to be $\mathrm{diam}(\xi):=\sup_{x\in\mathbb{T}^{3}}\mathrm{diam}(\xi(x))$. Fix a small $\varepsilon_{0}>0$ and let $\mathcal{Q}$ be the set of $\mu$-measurable partitions of diameter less than $\varepsilon_{0}$. For $\xi\in\mathcal{Q}$, we define a finer partition $\eta=:\mathcal{Q}^{u}(\xi)$ subordinated to the unstable foliation by defining $\eta(x)$ to be the connected component of $\xi(x)\cap\mathcal{W}^{u}(x)$ containing $x$ for any $x\in\mathbb{T}^{3}$. Let $\mathcal{Q}^{u}$ denote the set of all partitions obtained this way. The conditional entropy of $\xi$ given $\eta$ with respect to $\mu$ is given by
@@
where the value does not depend on $\xi$ or $\eta$, see [24, Theorem A].
@@
The topological counter part of the metric unstable entropy is the topological unstable entropy. It was introduced by Saghin and Xia [36]. It is equal to the exponential growth rate of the volume of strong unstable balls:
@@
where $\mathrm{Vol}$ is the induced volume by the Riemannian metric. The limit does not depend on either, $x\in\mathbb{T}^{3}$, $\delta>0$ or the choice of the Riemannian metric. Hu, Hua and Wu [24, Definition 1.4] gave an alternative definition of $h^{u}_{\mathrm{top}}(f)$ in terms of spanning or separated sets which we will not need in this paper. The link between the two notions of unstable entropies is given by the variational principle (1.1), see [24, Theorem D]. A measure $\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})$ is a measure of maximal unstable entropy (or $u$-MME) if $h^{u}_{\mathrm{top}}(f)=h^{u}_{\mu}(f)$. Their set is denoted by $\mathcal{M}^{u}(f)$. By the work of Hu, Wu and Zhu [25, Theorem B], the set $\mathcal{M}^{u}(f)$ is always non-empty, convex and compact. Moreover, its extreme points are exactly the ergodic $u$-MMEs.
@@ 2.3. Center Jacobian
where $\|.\|$ denotes the adapted Riemannian metric and $\alpha>0$. By the thermodynamical formalism (see [21, Theorem 20.3.7]), there is a unique equilibrium state $\mu_{J^{c}_{f}}$ associated to $J_{f}^{c}$. It is the unique invariant probability measure such that
@@
We use [24, Corollary A.1] which states that
@@
Now, the measure $\mu_{J^{c}_{f}}$ is ergodic so by the ergodic theorem and the chain rule, one has
@@ 3. Construction of a Margulis system of measures
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. In this section, we construct a system of $u$-measures for $f$. Recall that there exists a Hölder continuous homeomorphism $h:\mathbb{T}^{3}\to\mathbb{T}^{3}$ and a hyperbolic matrix $A\in\mathrm{GL}_{3}(\mathbb{Z})$ such that $h\circ f=A\circ h,$ see for instance [21, Theorem 18.6.1]. The matrix $A$ has three distinct and real eigenvalues $\lambda_{s}<1<\lambda_{c}<\lambda_{u}$ and the corresponding eigenlines are denoted by $E_{s}^{A}$, $E_{c}^{A}$ and $E_{u}^{A}$ respectively. We have $A\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and the corresponding foliations are given by $\mathcal{W}^{\bullet}_{A}(x)=x+E_{\bullet}^{A}$, for $\bullet=s,c,u$.
@@
    (3.2) $\|H-\mathrm{Id}\|_{\infty}:=\sup_{x\in\mathrm{R}^{3}}d(H(x),x)\leq L,$
@@
The first three facts were proven by Potrie in [34] and the last one was proven by Ren, Gan and Zhang in [42]. Note that the fact that $E_{s}$ and $E_{u}$ are not jointly integrable is an open and dense condition in $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, see [14, 7, 13].
We first show the following uniform volume bound.
@@
By Potrie [34, Corollary 7.7], the strong unstable foliation is quasi-isometric:
@@
We use a compactness argument similar to [9, 10].
@@
Note that $S$ leaves $\mathcal{X}$ invariant and is continuous. Since $\mathcal{X}$ is convex and compact, the Schauder-Tychonoff fix point Theorem implies that $S$ has a fix point $\mu^{u}$:
(3.13) $\exists\mu^{u}\in\mathcal{X}\quad f^{*}\mu^{u}=e^{h^{u}_{\mathrm{top}}(f)}\mu^{u}$
Since $\mu^{u}\in\mathcal{X}$, (3.4) implies that
(3.14) $\forall x\in\mathbb{T}^{3},\quad C_{1}(\delta)\leq\mu^{u}_{x}(\mathcal{W}^{u}(x,\delta))\leq C_{2}(\delta),$

### v3

@@ 2.1.1. Definition
are continuous and satisfy $\lambda_{x}^{s}<1<\lambda_{x}^{c}<\lambda_{x}^{u}$ for any $x\in\mathbb{T}^{3}$. In particular, $f$ is an Anosov map with stable bundle $E_{s}$ and unstable bundle $E_{cu}:=E_{c}\oplus E_{u}$. The bundles $E_{c}$ and $E_{u}$ are respectively called the center and strong unstable bundle of $f$. Finally, $E_{cs}:=E_{c}\oplus E_{s}$ is the center stable bundle of $f$. For a detailed introduction to Anosov diffeomorphisms, we refer to [undeft].
@@ 2.1.2. Invariant manifolds
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Since $f$ is Anosov, $E_{s}$ and $E_{cu}$ are integrable to $f$-invariant foliations $\mathcal{W}^{s}$ and $\mathcal{W}^{cu}$ called respectively the stable and unstable foliations of $f$, see [undeft, Theorem 6.2.3]. Moreover, since $E_{u}$ is the strong unstable foliation of a strongly partially hyperbolic diffeomorphism, it integrates to a $f$-invariant foliation $\mathcal{W}^{u}$ called the strong unstable foliation of $f$, see [undefaf, Theorem 4.1]. Actually, by the work of Brin, Burago and Ivanov [undefe] and Potrie [undefag], any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is dynamically coherent. This means that the center stable bundle $E_{cs}$ also integrates to a $f$-invariant foliation $\mathcal{W}^{cs}$ called the center stable foliation of $f$. The collection of leaves $\mathcal{W}^{c}(x):=\mathcal{W}^{cs}(x)\cap\mathcal{W}^{cu}(x)$ for $x\in\mathbb{T}^{3}$ defines the center foliation which integrates the center bundle $E_{c}$. For any $\bullet=s,c,u,cs,cu$, any $x\in\mathbb{T}^{3}$ and any $\delta>0$, let $\mathcal{W}^{\bullet}(x,\delta):=\{y\in\mathcal{W}^{\bullet}(x)\mid d^{\bullet}(x,y)<\delta\}$, where $d^{\bullet}$ is the distance on the leaf.
@@ 2.1.3. Local product structure and rectangles
Rectangles exist because of the local product structure, see [undefk, Eq. $(2.2)$].
@@ 2.1.4. Holonomies
The strong unstable foliation is absolutely continuous, see (5.5) for the exact statement we will need. This fact will be used in the proof of Proposition 5.2 to show the $\mathrm{Hol}^{u}$-invariance of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$. The center stable foliation is not always absolutely continuous and this explains why proving the $\mathrm{Hol}^{cs}$-invariance of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is harder. We refer to [undef, §2.2.3] for a more detailed discussion of the lack of regularity of center stable holonomies.
@@ 2.2.1. Conditional measures
We recall some facts about measurable partitions and conditional measures and refer to [undefu] for more details. Let $(X,\mu)$ be a probability space. For $\xi$ a partition of $X$ into $\mu$-measurable sets, we write $\xi(x)\in\xi$ for the element of the partition containing $x$. Let $\xi_{1},\xi_{2}$ be two partitions. We say that $\xi_{1}$ refines $\xi_{2}$ denoted $\xi_{2}\prec\xi_{1}$ if for $\mu$-a.e. $x\in X$, one has $\xi_{1}(x)\subset\xi_{2}(x)$. The joint of $\xi_{1}$ and $\xi_{2}$ is the partition defined by $\xi_{1}\vee\xi_{2}:=\{\xi_{1}(x)\cap\xi_{2}(x)\mid x\in X\}$. It is the least fine partition that refines both $\xi_{1}$ and $\xi_{2}$. A partition $\xi$ is said to be measurable if there is a sequence of finite partitions $\xi_{n}$ such that $\xi=\vee_{n=0}^{+\infty}\xi_{n}.$
@@
see [undefk, Equation $(2.6)$]. One defines the conditionals of $\mu$ on the strong unstable manifolds $(\mu^{u}_{x})_{x\in R}$ by replacing $cs$ by $u$ in the previous definition. Although this is not explicit in the notations, the conditionals depend on the rectangle $R$. Note however that the ambiguity only consists in a normalizing constant, see [undefk, Lemma 2.10].
@@ 2.2.2. Metric unstable entropy
We now recall the definition of the metric unstable entropy. For further details, we refer to [undefw]. In this subsection, we consider $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and $\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})$ a $f$-invariant probability measure. For a partition $\xi$, define its diameter to be $\mathrm{diam}(\xi):=\sup_{x\in\mathbb{T}^{3}}\mathrm{diam}(\xi(x))$. Fix a small $\varepsilon_{0}>0$ and let $\mathcal{Q}$ be the set of $\mu$-measurable partitions of diameter less than $\varepsilon_{0}$. For $\xi\in\mathcal{Q}$, we define a finer partition $\eta=:\mathcal{Q}^{u}(\xi)$ subordinated to the unstable foliation by defining $\eta(x)$ to be the connected component of $\xi(x)\cap\mathcal{W}^{u}(x)$ containing $x$, for any $x\in\mathbb{T}^{3}$. Let $\mathcal{Q}^{u}$ denote the set of all partitions obtained this way. The conditional entropy of $\xi$ given $\eta$ with respect to $\mu$ is given by
@@
where the value does not depend on $\xi$ or $\eta$, see [undefw, Theorem A].
@@
The topological counter part of the metric unstable entropy is the topological unstable entropy. It was introduced by Saghin and Xia [undefak]. It is equal to the exponential growth rate of the volume of strong unstable balls:
@@
where $\mathrm{Vol}$ is the induced volume by the Riemannian metric. The limit does not depend on either, $x\in\mathbb{T}^{3}$, $\delta>0$ or the choice of the Riemannian metric. Hu, Hua and Wu [undefw, Definition 1.4] gave an alternative definition of $h^{u}_{\mathrm{top}}(f)$ in terms of spanning or separated sets which we will not need in this paper. The link between the two notions of unstable entropies is given by the variational principle (1.1), see [undefw, Theorem D]. A measure $\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})$ is a measure of maximal unstable entropy (or $u$-MME) if $h^{u}_{\mathrm{top}}(f)=h^{u}_{\mu}(f)$. Their set is denoted by $\mathcal{M}^{u}(f)$. By the work of Hu, Wu and Zhu [undefx, Theorem B], the set $\mathcal{M}^{u}(f)$ is always non-empty, convex and compact. Moreover, its extreme points are exactly the ergodic $u$-MMEs.
@@ 2.3. Center Jacobian
where $\|.\|$ denotes the adapted Riemannian metric and $\alpha>0$. By the thermodynamical formalism (see [undeft, Theorem 20.3.7]), there is a unique equilibrium state $\mu_{J^{c}_{f}}$ associated to $J_{f}^{c}$. It is the unique invariant probability measure such that
@@
We use [undefw, Corollary A.1] which states that
@@
Now, the measure $\mu_{J^{c}_{f}}$ is ergodic so by the ergodic theorem and the chain rule,
@@ 3. Construction of a Margulis system of measures
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. In this section, we construct a system of $u$-measures for $f$. Recall that there exists a Hölder continuous homeomorphism $h:\mathbb{T}^{3}\to\mathbb{T}^{3}$ and a hyperbolic matrix $A\in\mathrm{GL}_{3}(\mathbb{Z})$ such that $h\circ f=A\circ h,$ see for instance [undeft, Theorem 18.6.1]. The matrix $A$ has three distinct and real eigenvalues $\lambda_{s}<1<\lambda_{c}<\lambda_{u}$ and the corresponding eigenlines are denoted by $E_{s}^{A}$, $E_{c}^{A}$ and $E_{u}^{A}$ respectively. We have $A\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and the corresponding foliations are given by $\mathcal{W}^{\bullet}_{A}(x)=x+E_{\bullet}^{A}$, for $\bullet=s,c,u$.
@@
    (3.2) $\|H-\mathrm{Id}\|_{\infty}:=\sup_{x\in\mathbb{R}^{3}}d(H(x),x)\leq L,$
@@
The first three facts were proven by Potrie in [undefag] and the last one was proven by Ren, Gan and Zhang in [undefaq]. Note that the fact that $E_{s}$ and $E_{u}$ are not jointly integrable is an open and dense condition in $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, see [undefm, undeff, undefl].
We first show the following uniform volume bounds.
@@
By Potrie [undefag, Corollary 7.7], the strong unstable foliation is quasi-isometric:
@@
We use a compactness argument similar to [undefh, undefi] and first construct a system of measure which has full support in each leaf and has the scaling property.
@@
Note that $S$ leaves $\mathcal{X}$ invariant and is continuous. Since $\mathcal{X}$ is convex and compact, the Schauder-Tychonoff fix point Theorem implies that $S$ has a fix point $\tilde{\mu}^{u}$:
(3.13) $\exists\tilde{\mu}^{u}\in\mathcal{X}\quad f^{*}\tilde{\mu}^{u}=e^{h^{u}_{\mathrm{top}}(f)}\tilde{\mu}^{u}$
Since $\tilde{\mu}^{u}\in\mathcal{X}$, (3.4) implies that
(3.14) $\forall x\in\mathbb{T}^{3},\quad C_{1}(\delta)\leq\tilde{\mu}^{u}_{x}(\mathcal{W}^{u}(x,\delta))\leq C_{2}(\delta),$